OpenStudy (anonymous):

(a+3)(a^3-2a^2+5a-2)

hero (hero):

lol. Suck your what?

OpenStudy (anonymous):

It's my lat name

hero (hero):

You have a strange name, lol

OpenStudy (anonymous):

multiply A+3 by everyterm in the other group i believe..than combine liketerms

OpenStudy (anonymous):

your last name is suckmylee??

OpenStudy (anonymous):

really??

OpenStudy (anonymous):

Lee

OpenStudy (anonymous):

so suck ur last name?

OpenStudy (anonymous):

and my initials S.M.L. now can someone remind me how to answer this?

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

so your first name is suck?

OpenStudy (anonymous):

Suck my replaces my real names Lee is my real last name

OpenStudy (anonymous):

(a+3)(a^3-2a^2+5a-2)= a(a^3-2a^2+5a-2)+3(a^3-2a^2+5a-2)= a^4-2a^3+5a^2-2a+3a^3-6a^2+15a-6= a^4+a^3-a^2+13a-6

hero (hero):

suckymylee, do you know what suck means?

OpenStudy (anonymous):

Thanks Jahtoday and yeah I do know -_- I think I'm going to change it now shesh